A water sample was run through the colorimetric procedure for the analysis of nitrate, giving 65.0% transmittance. A sample containing 1.00 ppm nitrate run through the exactly the same/identical procedure gave 24.6% transmittance. What was the concentration of nitrate in the first sample?


1
Expert's answer
2021-07-05T03:25:10-0400

The relation between absorbance and transmittance is given by A= 2-log (%T)

55% transmittance correspond to A= 2-log (55)= 0.2596

for 24.6% tranasmittance A= -log (24.6)= 0.6091

but absorance A= ebc

for A= 0.2596 = ebc ( e and b are constant s

A= 0.6091 =ebC1

C1/C= 0.6091/0.2596

C1=2.34 ppm

3. from A = ebc ( eb are constants)

0.075= ebC1 ( C1 is concentration of lead pure sample)

0.11 =ebC2

C2/C1=0.11/0.075=1.46

C2=6×1.46 =8.76 ug/L